3.Ryan is self-employed.  This year Ryan used his personal auto for several long business trips. Ryan paid $1,400 for gasoline on these trips. His depreciation on the car if he was using it fully for business purposes would be $3,000. During the year, he drove his car a total of 12,000 miles (combination of business and personal travel). Ryan can provide written documentation of the business purpose for trips totaling 3,000 miles. If Ryan does not use the standard mileage rate, what business expense amount can Ryan deduct (if any) for these trips?
Answer #1

The two expenses that Ryan can charge towards business expenses are (a) Gasoline expense, and (b) Depreciation expense


Total miles travelled = 12,000
Total miles pertain to business purpose = 3,000
therefore, percentage use of vehicle for business purpose = ( (3000 / 12000) * 100) = 25%
Therefore, the following expense can be allocated towards business:
(a) Gasoline expense = $1,400 * 25% = $350
(b) Depreciation expense = $3000 * 25% = $750

Total Business Expense amount can Ryan deduct for this trips = $1100

Hope this helps you, Thank you!?